2. A random variable X has this probability distribution:



x 0 1 2 3 4 5



P(X=x) 0.1 ? 0.05 0.1 0.03 0.3




Find f(2)




1
Expert's answer
2022-04-21T10:02:11-0400

"f(2)=P(X\u22642)=1-P(X>2)=1-P(X=3)-P(X=4)-P(X=5)=1-0.1-0.03-0.3=0.57"